### v5

# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

Abstract
Supply chain risk assessment (SCRA) is pivotal for ensuring resilience in increasingly complex global supply networks. While existing reviews have explored traditional methodologies, they often neglect emerging artificial intelligence (AI) and machine learning (ML) applications and mostly lack combined systematic and bibliometric analyses. This study addresses these gaps by integrating a systematic literature review with bibliometric analysis, examining 1,903 articles (2015–2025) from Google Scholar and Web of Science, with 54 studies selected through PRISMA guidelines. Our findings reveal that ML models, including Random Forest, XGBoost, and hybrid approaches, significantly enhance risk prediction accuracy and adaptability in post-pandemic contexts. The bibliometric analysis identifies key trends, influential authors, and institutional contributions, highlighting China and the United States as leading research hubs. Practical insights emphasize the integration of explainable AI (XAI) for transparent decision-making, real-time data utilization, and blockchain for traceability. The study underscores the necessity of dynamic strategies, interdisciplinary collaboration, and continuous model evaluation to address challenges such as data quality and interpretability. By synthesizing AI-driven methodologies with resilience frameworks, this review provides actionable guidance for optimizing supply chain risk management, fostering adaptability, and informing future research in evolving risk landscapes.
keywords:
Artificial Intelligence (AI) , Supply chain risk assessment , Bibliometric analysis , Content analysis , Systematic literature review
[…]
[inst3]organization=Department of Computer Science,addressline=American International University-Bangladesh (AIUB), city=Dhaka, postcode=1229, country=Bangladesh
1 Introduction
Supply chain management (SCM) has long grappled with various challenges from various sources. Past outbreaks of infectious diseases, geological upheavals like earthquakes, and other natural catastrophes have put supply chains (SCs) at risk, albeit on a limited scale (Govindan et al.,, 2020). These incidents have illustrated how interrelationships within SCs have formed intricate risk contagion networks, making them susceptible to contagion effects (Agca et al.,, 2022). The cascade effect, characterized by risk spillover from one enterprise to another, exacerbates the challenges, culminating in a chain reaction of SC Resilience (SCR) disasters (Roukny et al.,, 2018). Across the globe, companies confront formidable challenges across all stages of their SCs. Suppliers failing to meet delivery obligations, unpredictable shifts in customer demands, and episodes of panic buying are just a few examples of the hurdles companies face (Ivanov,, 2020). In pursuing market leadership, enterprises must also contend with the complexity introduced by the management and launch of innovation projects, further complicating SCM (Kwak et al.,, 2018). Furthermore, digital transformation has ushered in a wave of technological advancements in SC operations (Kwak et al.,, 2018).
However, the most significant disruption in recent times emerged during the COVID-19 pandemic, impacting global SCs profoundly. The disruptions encompassed not only the movement of people but also the flow of raw materials and finished goods, along with extensive disruptions in factory and SC operations (Sheng et al.,, 2021). These disruptions ushered in unprecedented challenges for SC professionals as they grappled with an entirely new reality (Araz et al.,, 2020; Craighead et al.,, 2020). The pandemic represents a distinctive form of SC disruption, distinct from natural or man-made disasters (e.g., the 2011 earthquake in Japan and the 9/11 attacks) and disruptions driven by evolving technologies and changing customer attitudes (Zhang et al.,, 2020). While existing guidance exists for predicting, managing, and responding to these disruptions, the challenges brought about by COVID-19 have underscored the paramount importance of risk management, unlike any previous disruption (Bode et al.,, 2011; Craighead et al.,, 2020). Consequently, the concept of SCR in COVID-19 has emerged as a critical area necessitating further exploration and development.
SCs are not the sole victims in this precarious environment, as the ripple effect extends to upstream and downstream enterprises. Practical risk management programs are imperative for enterprises to avert SCR and enhance SC risk management (SCRM). AI encompasses various techniques that enable autonomous decision-making in complex and uncertain environments. As noted by Baryannis et al., 2019b , an SCRM approach is considered artificially intelligent if it can autonomously decide on a course of action that leads to success in SCRM-related objectives while operating under a partially unknown SC environment. The AI techniques utilized in this context span from traditional symbolic AI, which relies on mathematical or knowledge-based problem representations, to sub-symbolic AI, including fuzzy systems and evolutionary computation, and statistical AI, which encompasses ML approaches. Among these, ML algorithms have showcased their ability to identify abnormal risk factors and derive predictive insights from historical data (Guo et al.,, 2021; Mohanty et al.,, 2021). By harnessing ML, enterprises can detect risk factors and anticipate market demands and potential risk scenarios (Punia et al.,, 2020; Wu et al., 2022a, ). ML’s proficiency in handling non-linear relationships further bolsters its superiority over traditional linear models. Additionally, its aptitude for processing unstructured data, a challenge where traditional models falter, positions ML as a formidable tool for addressing time, cost, and resource constraints within SCs. ML has emerged as the most suitable technique for risk assessment in SC. Therefore, this review focuses specifically on ML-based approaches in the context of SC resilience assessment.
However, despite its immense potential, SC researchers have historically exhibited limited familiarity with ML compared to other SCM aspects like mathematical programming and stochastic optimization (Liu et al.,, 2019; Shahed et al.,, 2021). Furthermore, the classification of ML algorithms remains unclear (Janiesch et al.,, 2021; Xu and Jackson,, 2019). Bridging this knowledge gap and exploring the value of ML for SCRM through interdisciplinary integration is a critical research need. Remarkably, there has been no prior effort to scrutinize the SCRM literature within an ML environment, in stark contrast to previous reviews centered on risk definition, classification, and management strategies. Companies should augment their analytical capabilities by harnessing organizational knowledge to enhance SCR, thereby elevating their information capabilities (Wong et al.,, 2020). As underscored by previous studies, the role of AI-enabled technologies extends to promoting innovations for enhanced SC performance (SCP) (Baryannis et al., 2019b, ; Nayal et al.,, 2021). The adaptation capabilities and information processing prowess offered by AI techniques hold the potential to enhance SCP (Belhadi et al., 2021a, ). Notably, AI has found application across diverse sectors, improving flexibility and communication and reducing undue fluctuations for successful project execution (Lalmi et al.,, 2021). To navigate the impact of risks and disruptions, (Katsaliaki et al.,, 2022) advocates integrating three critical facets: long-term partnerships, IT applications for business enhancement, and government policies that facilitate adaptability.
The contemporary SC landscape demands the integration of AI, specifically ML, to invigorate SCRA and SCRM. While traditional methods possess merit, their limitations can be effectively addressed by AI, which offers predictive capabilities, nonlinear relationship analysis, and unstructured data processing. As SCs continue to evolve, AI-based risk assessment is imperative for fostering resilience and sustaining the efficiency and effectiveness of SC operations. Existing reviews of SCRA using ML often fail to offer a comprehensive view of evolving ML techniques, neglecting emerging trends and overlooking diverse applications and non-journal sources. They also tend to overlook the role of ML in the response phases of SCRM and do not sufficiently address data-related challenges.
In light of the limitations associated with traditional SCRA methods, there is a growing interest in harnessing the potential of AI techniques to enhance risk assessment practices. This paper presents a systematic literature review (SLR) focused on the application of AI in SCRA. The existing literature lacks bibliometric analysis incorporating recent and adequate articles on the AI-SCRA topic. Another limitation is that most literature only contributed to bibliometric or content analysis. We tried to close this gap by including both analyses in our study, leveraging the systematic review methodology. The main objective of this review is to analyze the existing literature critically, identify research gaps, and provide insights into the use of AI techniques, such as ML, Deep Learning (DL), and natural language processing (NLP), to improve the accuracy and effectiveness of SCRA. Our paper has the following contributions-
  1. 1.
    Our study provides a comprehensive review of AI and ML applications in SCRA, supported by bibliometric and co-citation analyses to capture recent advancements and trends in the field.
  2. 2.
    We conduct content analysis to explore thematic trends, detailing the advantages, disadvantages, and specific application scenarios of various AI/ML models used in SCRA, offering insights into their suitability for different SC contexts.
  3. 3.
    Our proposed AI-based framework guides practitioners in assessing SC risks, from data collection to model selection and interpretation, facilitating efficient risk management in dynamic SCs.
The rest of the article is organized as follows: the “Literature Review” examines the current body of knowledge. The “Methodology” section details the study’s design, including search strategies and selection criteria. The “Bibliometric Analysis” provides insights from bibliometric tools, with visualizations and observations on AI in SCRA. A structured framework for SCRA is given in “Framework for SCRA using AI section. In the “Content Analysis of AI Techniques in SCRA,” the review discusses various AI techniques, highlighting their benefits, limitations, and real-world applications. Practical insights for the industry are provided in the “Practical and Managerial Implications” section, offering actionable guidance for practitioners. The review also outlines key “Challenges and Limitations” in this area, and suggests “Future Research Directions” for ongoing exploration. The paper concludes with a summary of findings and recommendations in the “Conclusion.”
2 Literature Review
Deiva Ganesh and Kalpana, (2022) conducted a comprehensive and descriptive literature study to identify AI and ML approaches in SCRM stages. This study examined SCRM research publications from three scientific databases from 2010 to 2021. They proposed a data analytics, simulation, and optimization framework that might produce a comprehensive SC risk identification, assessment, mitigation, and monitoring strategy. Hence, leveraging AI, Blockchain, and the Industrial Internet of Things (IIoT) to create smarter SCs can transform how firms handle uncertainty. This analysis of Deiva Ganesh and Kalpana, (2022) does not evaluate blockchain, big data, and IIoT-related SC articles, which may limit information. Due to limited research, only recent English-language publications on SCRM and AI were included. Nimmy et al., (2022) did a thorough literature evaluation on operational risk assessment methods and whether AI explains them. The report suggested that risk managers use auditable SC operational risk management methodologies to understand why they should take a risk management action rather than just what to do. While they employed the SLR method, they could only examine some AI solutions for SCRM. They suggested hybridizing SC operational risk management and explainable AI (XAI) to obtain XAI-like characteristics in future research. Li et al., (2023) analyzed COVID-19 SCR research history, present, and future. In particular, supervised ML classifies 1717 SCR papers into 11 subject categories. Each cluster was then studied in the context of COVID-19, indicating three related skills (interconnectedness, transformability, and sharing) on which enterprises could work to develop a more resilient SC post-COVID. Their data only came from Scopus’ Core Collection, which might affect the results; thus, they suggested adding WoS and EBSCOhost to the evaluation. (Xu et al.,, 2020). Given the fast expansion of SCR research, they only picked English-language articles, which may have excluded valuable knowledge. They recommended network analysis to determine cluster linkages and SCR literary themes.
Naz et al., (2022) examined the role of AI in building a robust and sustainable SC and offered optimal risk mitigation options. For review, 162 SCOPUS research publications were selected. Based on the nominated articles, Structural Topic Modeling produced various AI-related theme topics in SCR. AI research trends in SCR were examined using R-package bibliometric analysis. They solely studied journal articles, not conference papers, field reports, corporate reports, book chapters, etc. Ni et al., (2020) analyzed articles from 1998/01/01 to 2018/12/31 in five major databases to highlight the newest research trends in the field. ML applications in SCM were still developing due to a lack of high-yielding authors and poor publication rates. 10 ML algorithms were extensively utilized in SCM, but their utilization was unequal among the SCM tasks most often reported. This paper has limitations in reviewing only five popular databases to limit articles for evaluation that might have filtered some related articles. Second, only widely used ML methods in SCM were counted in this review, and other ML techniques brought to SCM may be helpful later on. Low-frequency ML algorithms should be analyzed for further research in this area. Finally, their article contained 32 well-recognized ML methods; however, some newly generated ML algorithms may be used in SCM after 2015. Baryannis et al., 2019b examined SC risk definitions and uncertainty. Then, a mapping analysis categorizes available literature by AI approaches and SCRM tasks. Most of the examined works focus on building and assessing a mathematical model that accounts for various uncertainties and hazards but less on establishing and analyzing the applicability of the suggested models. They found that only 9 of 276 research (3%) use comprehensive techniques that cover all three SCRM stages (identification, assessment, and response). Yang et al., (2023) thoroughly examined the advancement of ML algorithms in SCRM by gathering 67 publications from 9 authoritative databases in the first half of 2021. They analyzed only English language journal articles from 9 relevant academic databases, excluding conference papers, textbooks, and unpublished articles and notes. This analysis relied on keyword searches, which may have missed some work. Schroeder and Lodemann, (2021) used a comprehensive and multi-vocal literature study to get a complete picture of our subject field. The SLR identified 533 papers in this area and examined 23. The comprehensive literature study found that just a few examples of ML in SCRM have been reported in depth in the scientific literature, and those examples focus on manufacturing, transport, and the complete SC. Shi et al., (2022) examined 76 works on credit risk utilizing statistical, ML, and DL methods over the last eight years. They offered a unique classification approach and performance rating for ML-driven credit risk algorithms utilizing public datasets. Data imbalance, dataset inconsistency, model transparency, and DL model underuse are discussed. Their review found that most DL models outperform standard ML and statistical algorithms in credit risk prediction, and ensemble techniques outperform single models.
Previous studies in SCRM have shown limitations in evaluating certain technologies, such as blockchain, big data, and the IIoT, potentially limiting the scope of analysis. Our review addresses this gap by incorporating an analysis of these technologies within the context of SCRM, providing a more comprehensive assessment. There is a recognized need in the literature to integrate SC operational risk management with XAI for future research. Our review contributes to closing this gap by exploring the integration of these approaches and discussing their implications for SCRM. Some studies have highlighted the importance of utilizing multiple databases to evaluate SCRM literature comprehensively. Our review extends this by considering insights from the most reliable and prominent two databases, enhancing the breadth of our analysis. Previous research has found that only a small percentage of studies cover all three stages of SCRM comprehensively. Our review contributes to closing this gap by providing a holistic assessment of SCRM stages, including identification, assessment, and response strategies, leveraging insights from diverse studies.
3 Methodology
3.1 Study Design
The research utilized a systematic approach to conduct a comprehensive literature review and a bibliometric analysis. This methodology aimed to scrutinize the intersection of AI/ML techniques and SCRA, providing insights into research trends, methodologies, and emerging themes. We formulated the following research questions to guide our SLR:
  • RQ1:
    What is the current state of research on applying AI/ML techniques in SCRM?
  • RQ2:
    Which AI/ML techniques are commonly employed in SCRA, prediction, and mitigation?
  • RQ3:
    What are the key findings and trends identified in the literature regarding using AI/ML in SCRM?
  • RQ4:
    What are the research gaps and future directions in this field?
Our approach is guided by the structured methodology of Denyer and Tranfield, (2009) for systematic literature reviews, ensuring rigor in study identification and synthesis. Additionally, we adhered to the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework to enhance transparency and minimize bias in the study selection process. This approach aimed to identify and analyze research articles on integrating AI/ML techniques within SCRA.
3.2 Search Strategy
We conducted a comprehensive search using two major academic databases: Google Scholar and Web of Science (WoS) Core Collection. The WoS Core Collection, recognized as a high-quality digital literature database, has gained widespread acceptance among researchers and is currently regarded as one of the most appropriate databases for bibliometric analysis (Wang et al.,, 2024). Compared to other databases, WoS offers more extensive data coverage and provides data formats compatible with analysis tools such as VOSviewer. We used the “Publish or Perish 8” (Harzing,, 2016) software for the Google Scholar search to extract all data of the most relevant articles suitable for bibliometric analysis. The initial search on Google Scholar using Publish or Perish yielded 1000 articles, and WoS, using advanced search, yielded 903 articles before specifying any date range. We employed the following search strings to retrieve relevant articles:
(ALL=( “supply chain”) AND (“risk*” OR “credit risk*” OR “risk assessment*” OR “risk prediction” OR “disruption*” OR “disease outbreak*” OR “post COVID resilienc*”) AND (“artificial intelligence” OR “machine learning” OR “neural network*” OR “deep learning” OR “reinforcement learning” OR “SVM” OR “support vector machine” OR “boosting” OR “ensemble” OR “bayesian network model” OR “random forest” OR “LSTM” OR “long short-term memory”))) NOT (DT==(“PROCEEDINGS PAPER” OR “BOOK CHAPTER” OR “EDITORIAL MATERIAL” OR “LETTER”))
Table 1: Inclusion and exclusion criteria and justifications
Criterion Inclusion/Exclusion Justification
Date range 2015–2025 Focuses on recent studies, ensuring relevance to current advancements in AI for SCRA.
Language English only Ensures consistency in analysis and accessibility for an English-speaking academic audience.
Duplicate removal Exclusion Avoids redundancy and ensures each study contributes uniquely to the analysis.
Publisher reliability Exclusion Maintains academic quality by including only studies from reputable, peer-reviewed journals.
Publication type Exclusion of Q3/Q4, proceedings, book chapters, etc. Focuses on high-impact research, prioritizing quality over quantity.
Checking for relevance (title, abstract, and conclusion) Exclusion Ensures alignment with the study’s objectives based on content relevant to SC risk and AI.
Unrelated methodologies Exclusion Removes studies that employ computational models or frameworks that do not align with the research scope.
Limited generalizability Exclusion Excludes studies focusing on highly specific contexts that cannot be applied to broader AI-driven SCRA applications.
3.3 Article Selection Process
To ensure the selection of high-quality and relevant studies, we followed a systematic, multi-stage screening process guided by the PRISMA framework. The selection process was designed to minimize potential bias and ensure transparency at each stage. The stages are detailed as follows:
  1. 1.
    Date Range Filtering: We restricted our review to studies published between 2015–2025, excluding 23 articles, ensuring the inclusion of recent and relevant research. This step ensures that each study contributes uniquely to the analysis and aligns with the latest advancements in the field.
  2. 2.
    Removal of Duplicates: To avoid redundant analysis, we first identified and removed duplicate articles between the two databases, excluding 451 duplicate entries. This step is standard in systematic reviews to ensure each study contributes uniquely to the analysis.
  3. 3.
    Filtering by Publisher Reliability: We excluded 398 articles published by sources identified as less reliable, focusing only on those published in reputable publishers (IEEE, Springer Nature, Elsevier, Taylor & Francis, MDPI, Sage, Wiley, Nature Portfolio, ACM, etc.). This criterion helps maintain high academic integrity and ensures the final selection is based on studies that meet established research quality standards.
  4. 4.
    Non-English Exclusion: Since this review focuses on an English-speaking academic audience, we excluded 3 articles not written in English. This ensures that the included studies are accessible and can be consistently interpreted without translation bias.
  5. 5.
    Filtering by Publication Type: To prioritize peer-reviewed journal articles, we excluded papers classified as Q3 or Q4 and proceedings papers, book chapters, editorials, and letters, removing 244 articles. This criterion ensures that our review is based on high-impact research published in top-quality journals, as lower-ranked journals or other types of publications may not always meet rigorous peer-review standards.
  6. 6.
    Checking for Relevance (Title, Abstract, and Conclusion Screening): The remaining articles were reviewed based on their title, abstract, and conclusions to assess their relevance to the research questions of this study. Articles that did not focus specifically on AI applications in SCRA were excluded. This step eliminated 515 articles.
  7. 7.
    Exclusion of Studies with Unrelated Methodologies: After relevance screening, we further analyzed the methodologies used in the remaining studies. Articles that employed computational models or frameworks that did not align with our scope were removed. This step excluded 128 articles, ensuring the final selection focused on studies directly relevant to the research objectives.
  8. 8.
    Exclusion of Studies with Limited Generalizability: We excluded 81 highly context-specific studies that lacked generalizability to broader AI-driven SCRA applications, leaving a final sample of 54 articles for in-depth review. These studies primarily focused on niche use cases that do not translate well to wider SCRA applications.
Each criterion was carefully chosen to ensure that only studies of high relevance and quality were included, thus minimizing potential biases. Using dual databases (Google Scholar and WoS) ensured comprehensive coverage, while the sequential exclusion criteria systematically filtered out irrelevant or low-quality studies. To demonstrate further transparency, Table 1 outlines the inclusion and exclusion criteria and their justifications.
3.4 Bias Mitigation
To address concerns about potential bias, we followed a systematic approach that minimizes subjective selection at each stage:
  1. 1.
    By using two widely recognized databases, Google Scholar and WoS, we ensured broad initial coverage of the literature on AI in SCRA, reducing the likelihood of selection bias by capturing a diverse range of studies.
  2. 2.
    Each exclusion criterion was pre-defined and systematically applied, following a step-by-step approach that minimized subjective judgment. This ensures that all articles are assessed based on consistent standards.
  3. 3.
    The PRISMA framework was applied rigorously to eliminate irrelevant studies at multiple stages, further reducing the risk of bias by ensuring that only studies meeting high-quality and relevance standards were retained.
3.5 Final Selection
After applying all criteria, 54 articles were deemed relevant and of high quality for inclusion in this review. The selected articles cover various AI methodologies applied to different SC risk scenarios, providing a comprehensive foundation for our analysis. According to Figure 1, 778 articles were included in the bibliometric analysis to provide a broad understanding of the field, considering various quantitative metrics. The 54 shortlisted articles represent a more focused subset. We selected them based on additional qualitative criteria, such as relevance, methodology, and contribution to the research.
Figure 1: Prisma flow diagram for systematic literature review.
3.6 Data Extraction and Synthesis
The selected articles underwent detailed analysis, focusing on methodologies, AI/ML techniques employed, key findings, and implications for SCRA. Data extraction forms were filled in to gather pertinent information for synthesis and further analysis. To understand the research landscape comprehensively, we performed bibliometric and cocitation analysis of the 778 articles resulting from the filtering process. The bibliometric analysis helped us identify the field’s most influential authors, journals, and key research themes. Cocitation analysis allowed us to identify clusters of related articles and determine their interconnections. Our study has several limitations that should be acknowledged. First, the search was limited to articles published between 2015 and 2025, and it is possible that relevant articles published before or after this timeframe were not included.
4 Bibliometric Analysis
To comprehensively analyze the scholarly landscape of SCRA using AI, we employed a combination of tools, including MS Excel and VOSviewer. This bibliometric analysis encompasses several dimensions, offering a deeper understanding of the field by identifying trends, contributions, and intellectual structures.
4.1 Publication Trend Over Time
Figure 2: Annual publication trend (2015-2025) of the 778 published literature on AI in SCRA topic.
Figure 2 illustrates the annual publication trends from 2015 to 2025. A significant rise in publications is observed post-2020, likely influenced by the COVID-19 pandemic, highlighting the critical need for resilient SC systems. The peak in research output during 2024 reflects the heightened importance of AI/ML-driven risk assessment in tackling complex SC challenges. However, the slight decline in 2025 is because of the first quarter of the year, which is expected to peak following the exponential trend line.
4.2 Geographical Distribution of Research
Figure 3: Geographical distribution of the authors affiliated with 778 published papers on AI in SCRA topic.
Figure 3 visualizes the geographical distribution of research contributions. China leads with 39.49% of 778 publications, followed by the United States (11.72%) and England (11.59%). This dominance underscores the strong focus of Chinese institutions on AI-driven SCRA. The global spread of contributions emphasizes the universal relevance of SCRA, with increasing participation from emerging economies.
Table 2: Top 10 authors according to the published papers
{tblr}
column2 = c, hline1-2,12 = -, Authors & Published Paper Count
Gupta S 14
Liu YK 12
Brintrup A 11
Kumar A 11
Mangla SK 8
Modgil S 8
Choi TM 7
Liu Y 7
Li J 6
Liu C 6
4.3 Analysis of Authors
As shown in Table 2, the top 10 authors have significantly influenced the domain of SCRA using AI/ML. Gupta S leads the field with 1.78% of the 778 publications, followed by Liu YK (1.53%) and Brintrup A (1.40%), who have contributed substantially to advancing AI-driven risk assessment. Other key contributors include Kumar A (1.40%), Mangla SK (1.02%), and Modgil S (1.02%), each demonstrating a strong research footprint in this domain. Additionally, Choi TM (0.89%), Liu Y (0.89%), Li J (0.76%), and Liu C (0.76%) round out the top 10 authors, further emphasizing the interdisciplinary and global nature of research in AI-based SCRA. These authors have played a crucial role in developing AI-driven risk assessment frameworks, predictive models, and decision-support systems, shaping the future of SC resilience and risk mitigation strategies.
4.4 Publisher Contributions
Figure 4: Publisher-wise distribution of 778 articles.
Figure 4 presents the contributions of various publishers. Elsevier is the most prominent, contributing over 262 articles, followed by Springer Nature, MDPI, Taylor & Francis, and IEEE. These publishers serve as key platforms for disseminating high-quality research, reflecting their alignment with the evolving needs of AI and SCRA scholarship.
(a) Author co-citations network graph.
(b) Co-citation network graph of cited references.
(c) Co-citation network graph of publication sources.
Figure 5: Co-citation network graphs visualization using (a) authors, (b) cited references, and (c) sources as the unit of analysis.
4.5 Co-citation Analysis
Co-citation analysis was conducted to explore the intellectual structure of AI-SCRA research, focusing on authors, references, and publication sources. Figure 5(a) presents an author co-citation network comprising 55 authors grouped into three distinct clusters. The visualization highlights key contributors to AI-driven SCRA, with Ivanov and Baryannis playing central roles in the field. Other notable researchers, such as Dubey, Choi, and Zhu, also exhibit strong influence within their respective clusters. Figure 5(b) presents a co-citation network with three main clusters, highlighting key contributions to AI-driven SCRM. Notable references, including Ivanov (2019, 2020), Baryannis (2019), and Tang (2006), emphasize hybrid AI models, risk prediction frameworks, and post-COVID resilience strategies. Lastly, Figure 5(c) underscores the centrality of academic journals, with the International Journal of Production Research standing out as a key publication venue, reflecting its substantial impact on the field’s scholarly discourse and research progression.
(a) Co-authorship network analysis of 48 authors.
(b) Co-authorship network analysis of 18 countries.
Figure 6: Co-authorship network graphs visualizing collaborations at (a) author and (b) country levels.
4.6 Co-authorship Analysis
Co-authorship analysis is a powerful method for understanding collaborations among authors and researchers based on joint publication activities. This analysis highlights the collaborative nature of the research domain, uncovering influential partnerships and the significant contributions of specific countries. Figure 6(a) illustrates the co-authorship network among authors in this field. Using a threshold of at least one document and a minimum of 55 citations per author, 48 authors out of a total of 304 were identified as significant contributors. The visualization, generated using VOSviewer, reveals four distinct clusters, color-coded as red, green, blue, and yellow. These clusters represent groups of authors with strong collaborative ties. For example, the robust connections surrounding “Modgil” in the red cluster indicate frequent co-authorship and shared research initiatives, reflecting a high degree of collaboration within the cluster. Figure 6(b) showcases the co-authorship network at the country level. By applying thresholds of at least two documents and nine citations per country, 18 out of 36 countries were included in the analysis. The visualization highlights China, USA, and England as leading contributors, dominating document count, citation impact, and total link strength. This dominance underscores their pivotal roles in advancing AI/ML-driven SCRA research. The results also emphasize strong international collaborations between these countries and other key players, fostering a globally interconnected research ecosystem.
(a) Co-occurrence network analysis of top 26 authors’ keywords.
(b) Co-occurrence network analysis of top 47 keywords used in the literature related to SCRA-AI.
Figure 7: The Co-occurrence Network visualizes the intricate relationships among identified keywords. Each node represents a keyword, and the edges denote the frequency of co-occurrence. Clusters of tightly connected nodes signify thematic groupings, providing a comprehensive snapshot of the intellectual structure within the analyzed literature.
4.7 Co-occurrence Analysis
The co-occurrence analysis offers valuable insights into the thematic structure of SCRA and AI/ML research by identifying how keywords are interconnected within the scholarly literature. Using VOSviewer, we conducted this analysis by focusing on the units of analysis, such as “Authors’ Keywords” and “Keywords Plus.” This approach reveals latent connections and thematic clusters, enabling a deeper understanding of this domain’s critical topics and prevalent themes. The network map shown in Figure 7(a) illustrates the relationships among the 26 most frequently occurring author keywords. These keywords were selected from an initial pool of 1,483, applying a minimum threshold of three occurrences. Larger nodes, such as “artificial intelligence” and “machine learning,” indicate higher frequencies. The proximity of the nodes reflects their degree of association, with closely grouped nodes forming thematic clusters. Four clusters emerged, highlighting key themes. Figure 7(b) illustrates the co-occurrence of keywords in the reviewed literature. The analysis identified 47 frequently occurring terms from 1,032, using a minimum threshold of three occurrences. These keywords were categorized into five distinct clusters, reflecting prevalent themes and connections within the research field. Terms such as “management” and “framework” exhibited the strongest linkages, indicating their crucial roles in connecting different thematic areas. Additionally, other significant clusters highlighted emerging trends, including “supply chain resilience”, “big data”, “blockchain” and “sustainability”. This co-occurrence analysis offers a comprehensive overview of the intellectual structure of SCRA research, showcasing established and emerging themes. The findings emphasize the interconnectivity of concepts and provide guidance for future research directions by identifying critical areas of scholarly focus.
Table 3: Top 10 Affiliations According to Paper Count
Affiliations Published paper count % of 778 articles
HONG KONG POLYTECHNIC UNIVERSITY 22 2.799
CHINESE ACADEMY OF SCIENCES 21 2.672
INDIAN INSTITUTE OF TECHNOLOGY (IIT) SYSTEM 20 2.545
UNIVERSITY OF TEHRAN 19 2.417
INDIAN INSTITUTE OF MANAGEMENT (IIM) SYSTEM 15 1.908
NEOMA BUSINESS SCH 15 1.908
UNIVERSITY OF CAMBRIDGE 15 1.908
HEBEI UNIVERSITY 14 1.781
EMLYON BUSINESS SCHOOL 11 1.399
INDIAN INSTITUTE OF TECHNOLOGY (IIT) DELHI 11 1.399
Others 615 79.26
4.8 Top 10 Affiliations
Table 3 highlights the top 10 affiliations contributing to SCRA research, ranked by their published paper count. Hong Kong Polytechnic University leads the list with 22 publications (2.80%), followed closely by the Chinese Academy of Sciences (2.67%) and the Indian Institute of Technology (IIT) System (2.55%). Other key contributors include the University of Tehran (2.42%), the Indian Institute of Management (IIM) System (1.91%), and NEOMA Business School (1.91%), all of which have played a crucial role in advancing AI-driven SCRA methodologies. Prominent European institutions such as the University of Cambridge (1.91%), Hebei University (1.78%), and Emlyon Business School (1.40%) further highlight the global nature of SCRA research, emphasizing interdisciplinary contributions across Asia, Europe, and the Middle East. Notably, IIT Delhi (1.40%) is a significant research hub in this field. Together, these top 10 affiliations account for a substantial share of high-impact studies in SCRA, driving advancements in AI applications for risk mitigation. Meanwhile, the “Others” category, comprising 615 institutions, represents 79.26% of the contributions, reflecting the diverse and widespread nature of research in this domain.
4.9 Bibliographic Coupling
Bibliographic coupling assesses the similarity of documents, authors, or organizations based on shared references, revealing thematic alignments and intellectual structures. Key insights are as follows:
4.9.1 Authors
Figure 8 illustrates the bibliographic coupling analysis of 28 authors, grouping them into five thematic clusters based on shared references, highlighting intellectual connections in SCR research. Prominent contributors like Lim M.K., Gupta, Shivam, Wang G.-J., and Hussain F.K. demonstrate centrality, reflecting their significant influence in areas such as risk prediction, disruption management, and quantitative modeling. The clusters reveal interdisciplinary themes, including resilience strategies, industry-specific applications, and the integration of AI-driven methods. This analysis showcases established research directions and emerging trends, providing a roadmap for future studies and collaborations in SCR.
Figure 8: Bibliographic coupling analysis of 28 authors.
4.9.2 Countries
Figure 9 illustrates the bibliographic coupling among 18 countries, categorizing them into clusters highlighting global research collaborations in AI-driven SCRA. Countries such as China, the United States, and England emerge as central hubs, showcasing their pivotal roles in advancing research and promoting international partnerships. The clusters signify thematic alignments across regions, with nations collaborating on critical topics like predictive analytics, sustainability, and resilience strategies.
Figure 9: Bibliographic coupling analysis of 18 countries.
4.9.3 Documents
Figure 10 illustrates the bibliographic coupling analysis of 62 documents, visually depicting the thematic clustering and citation relationships among the studies. The eight clusters highlight distinct research themes within the field, showcasing the interconnectedness of key works. Highly influential documents, such as those by “Baryannis G.”, “Cavalcante I. M.”, and “Zhu Q.”, emerge as central nodes, reflecting their significant contributions to the discourse on SCRA using AI. This visualization underscores the collaborative and evolving nature of the research domain, providing insights into how various studies align and contribute to specific thematic advancements.
Figure 10: Bibliographic coupling analysis of 62 documents.
4.9.4 Organizations
Figure 11 presents the bibliographic coupling analysis of 24 organizations, highlighting their collaborative and thematic alignments in AI-based SCRA research. The visualization categorizes institutions into three clusters, highlighting key contributors such as the Hong Kong Polytechnic University, the University of Cambridge, and Nanjing University of Posts and Telecommunications. This analysis underscores the global distribution of research efforts and the leadership roles of prominent institutions in advancing AI-driven methodologies for SCRA.
Figure 11: Bibliographic coupling analysis of 24 organizations.
4.9.5 Journals
Figure 12 illustrates the bibliographic coupling analysis of the core academic journals driving the advancements in SCRA research by showing the interconnections and thematic clustering among 17 influential journals. Journals such as the “International Journal of Production Research” and “Annals of Operations Research” play a crucial role, providing foundational studies and methodologies that inform AI-driven approaches to SCRA. These clusters reveal thematic trends and emerging focus areas and facilitate knowledge dissemination by connecting researchers across disciplines.
Figure 12: Bibliographic coupling analysis of 17 publications sources.
Table 4: Publication Trends by Journal (2015-2025)
Source Title Publication Year Grand Total
2015 2016 2017 2018 2019 2020 2021 2022 2023 2024 2025
INTERNATIONAL JOURNAL OF PRODUCTION RESEARCH 1 5 3 4 18 7 38
SUSTAINABILITY 1 2 4 5 9 9 3 1 34
ANNALS OF OPERATIONS RESEARCH 2 10 18 1 31
ENGINEERING APPLICATIONS OF ARTIFICIAL INTELLIGENCE 1 1 2 2 5 8 2 21
INTERNATIONAL JOURNAL OF PRODUCTION ECONOMICS 1 1 1 1 1 6 8 2 21
COMPUTERS & INDUSTRIAL ENGINEERING 2 1 3 2 3 5 4 20
IEEE ACCESS 2 4 1 2 4 3 16
IEEE TRANSACTIONS ON ENGINEERING MANAGEMENT 4 12 16
EXPERT SYSTEMS WITH APPLICATIONS 1 1 1 2 3 4 3 15
TRANSPORTATION RESEARCH PART E-LOGISTICS AND TRANSPORTATION REVIEW 1 1 1 3 1 4 11
JOURNAL OF CLEANER PRODUCTION 2 5 1 2 10
APPLIED SCIENCES-BASEL 1 2 1 5 9
INDUSTRIAL MANAGEMENT & DATA SYSTEMS 1 3 1 2 7
JOURNAL OF INTELLIGENT & FUZZY SYSTEMS 1 2 3 1 7
PRODUCTION AND OPERATIONS MANAGEMENT 4 1 2 7
TECHNOLOGICAL FORECASTING AND SOCIAL CHANGE 1 2 2 2 7
APPLIED SOFT COMPUTING 1 1 1 1 2 6
DECISION SUPPORT SYSTEMS 1 3 2 6
EUROPEAN JOURNAL OF OPERATIONAL RESEARCH 1 1 2 1 1 6
INDUSTRIAL MARKETING MANAGEMENT 1 1 1 3 6
INTERNATIONAL JOURNAL OF LOGISTICS-RESEARCH AND APPLICATIONS 3 3 6
INTERNATIONAL JOURNAL OF PHYSICAL DISTRIBUTION & LOGISTICS MANAGEMENT 1 2 1 2 6
OPERATIONS MANAGEMENT RESEARCH 3 1 2 6
PLOS ONE 1 1 1 2 1 6
WIRELESS COMMUNICATIONS & MOBILE COMPUTING 3 3 6
Others 2 2 4 9 27 28 46 77 68 151 40 454
Grand Total 4 5 5 13 39 48 80 124 132 256 72 778
4.10 Journal Publication Trends
Table 4 presents a comprehensive breakdown of published papers in various source titles from 2015 to 2025. These source titles encompass a range of journals and publications contributing to the field of SCRA. The table clearly shows the publication trends, indicating how many papers were published in each source title each year and their cumulative total. It provides valuable insights into the distribution of research output and the source titles that have been active in this area.
5 Framework for SCRA using AI
Figure 13: Framework for SCRA using AI.
Our proposed framework for AI-driven SCRA provides a comprehensive pipeline for SC professionals to integrate AI techniques effectively. Figure 13 visually shows the sequential steps involved, from defining key risk indicators to implementing preventive measures. Following these steps, practitioners can systematically define risk indicators, source and pre-process data, select and train models, and interpret results to facilitate informed decision-making. The framework is enriched with insights from existing studies on AI and ML applications in SC risk, presenting best practices, key considerations, and real-world challenges and opportunities identified in the literature.
5.1 Defining Key Risk Indicators
Our first step is identifying the key input and output variables that drive SCRA. Input variables include transactional data, supplier metrics, and external factors like political or environmental conditions, while output variables classify risks into severity levels such as “low,” “medium,” and “high” to align with organizational decision-making priorities. Defining these variables is crucial, as they represent the foundation upon which AI models assess risk. Wu, (2024) emphasize capturing operational and risk-related data for comprehensive risk assessments, with Xu et al., (2024) using historical sales and market trends for disruption prediction. Including financial and non-financial indicators, as highlighted by Wang et al., (2021), enhances the assessment of risk exposure. Practitioners should ensure these variables align with their industry’s risk characteristics to produce actionable and context-relevant insights.
5.2 Data Collection
Effective risk assessment relies on internal and external data sources to capture a comprehensive picture of potential risks. Internal data from ERP systems, procurement records, and inventory management systems provides insights into operational processes. External data, such as market indicators, social media sentiment, and climate reports, are critical for Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. ERP systems, supplier systems, and logistics systems, alongside external sources like social media and weather forecasts. Such data covers both structured (e.g., numerical data) and unstructured formats (e.g., text or multimedia) to enhance risk detection. Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. The author used Twitter streaming API as a primary source for data extraction. Kosasih et al., (2022) utilized external platforms like Marklines for automotive SC data and Achilles Information Ltd for energy sector insights. Pan and Miao, (2023) collected data through surveys from industry practitioners to capture qualitative and contextual insights about SCR.Including such data enhances the model’s ability to preemptively detect risks like demand fluctuations or reputational issues, underscoring the value of a multimodal data approach.
5.3 Data Preprocessing
Preprocessing is a vital step. Raw data must undergo preprocessing to enhance model accuracy and efficiency. This step involves cleaning data to remove noise, normalizing values, and encoding categorical variables. Dimensionality reduction techniques like principal component analysis (PCA) applied by Luo et al., (2022) and imputation methods like residual mean replacement applied by Zhang et al., (2021) address challenges such as high-dimensional datasets and missing values. Time-series data, common in SC applications, requires additional adjustments like seasonality corrections. Techniques like Isolation Forest and Local Outlier Factor (LOF) detect anomalies, enhancing prediction reliability. SPSS statistical preprocessing methods have also been employed effectively in by Dang et al., (2022).
5.4 Model Selection
Practitioners have a range of model choices, each suited to different data characteristics and objectives in SCRA. Classical models like Support Vector Machine (SVM), Decision Tree (DT), and Logistic Regression (LR) are ideal for well-structured, low-dimensional datasets, where the data relationships are more straightforward and computational efficiency is prioritized. DL models, such as Convolutional Neural Networks (CNNs) and Long Short-Term Memory (LSTM) networks, are valuable for handling complex, high-dimensional data, including time-series demand data, where capturing temporal dependencies is essential. Hybrid models, which combine conventional ML and DL approaches, balance interpretability and accuracy, catering to scenarios where insights into model decisions are as important as predictive performance. Ensemble methods like Random Forest (RF) and XGBoost aggregate outputs from multiple models to enhance predictive accuracy and robustness. Additional research emphasizes that model selection should consider data availability, computational resources, and the necessity for model interpretability. Highly interpretable models are particularly important in regulated industries, such as finance, where transparency is critical. Hybrid and ensemble models, increasingly popular in recent literature, often effectively compromise predictive power and interpretability, making them suitable for complex SC risk scenarios.
5.5 Model Training and Hyperparameter Tuning
Model training typically involves cross-validation to improve model robustness. Tuning parameters like learning rate, depth, and number of estimators (in ensemble methods) is crucial to optimize performance. Tailored hyperparameter tuning can significantly enhance model performance in SC applications. Many researchers highlight the importance of automated tuning techniques, such as grid search, random search, and Bayesian optimization, to streamline model refinement and ensure that the final model is well-suited for specific risk scenarios. Wu et al., 2022b emphasized iterative optimization, while Zhang et al., (2022) addressed data imbalance with SMOTENC. Practitioners may also consider transfer learning approaches to leverage knowledge from similar datasets, especially in situations with limited labeled data.
5.6 Prediction
Once trained, the models generate predictions in the form of classifications (e.g., “high risk” vs. “low risk”), severity scores, or probabilities, which guide risk mitigation strategies. Baryannis et al., 2019b highlighted the role of predictive analytics in shielding SCs from disruptions. These predictions should ideally integrate into real-time dashboards to allow continuous monitoring and provide quick insights into risk fluctuations. This step facilitates proactive risk management, enabling organizations to adjust strategies promptly based on predicted risk levels and preventing disruptions before they escalate.
5.7 Interpretation of Results
Interpretability makes AI-driven predictions actionable, especially for stakeholders without a technical background. XAI, such as SHAP values or LIME, helps practitioners understand which variables most influence predictions. According to Olan et al., (2024), techniques like SHAP and LIME enhance decision-making by elucidating the rationale behind ML predictions, enabling better interpretability and fostering trust in SCRA. Interpretation tools also validate results with domain experts, ensuring the model’s output aligns with real-world SC conditions and providing a basis for actionable policy recommendations.
5.8 Preventive Measure
Based on the risk predictions and insights, practitioners can establish preventive strategies to mitigate potential SC disruptions. Common actions include adjusting inventory buffers, diversifying suppliers, or modifying production schedules. Preventive measures should be continuously refined based on AI predictions and risk evaluations. Incorporating feedback loops that allow practitioners to assess the impact of preventive measures on risk reduction is a good practice. Such loops can help organizations refine their approach over time, ensuring that the AI-driven risk assessment framework remains adaptable to changing conditions.
Table 5: Methods used and the key findings from the review articles.
Reference(s) Methods/Models Key Findings
Katsaliaki et al., (2022) Tensorization-based neural symbolic model Outperformed SNLP and GNN-LP in predicting relationships (AUC: 0.82, AP: 0.72).
Wang et al., (2022) Cost-sensitive learning to a RF (CSL-RF) model Improved prediction accuracy (Accuracy: 97.22%, AUC: 0.9850).
Wang et al., (2021) Adaptive heterogeneous multiview graph learning Improved forecasting accuracy (Accuracy: 87.44%).
Jianying et al., (2021) Optimized Backpropagation Neural Network (BPNN) and Particle Swarm Optimization-Backpropagation (PSO-BP) Outperformed the single Backpropagation (BP) (MAE: 2.1022, $R^{2}$: 0.932).
Chen et al., (2021) LSTM for forecasting LSTM model suitable for risk prediction (MAE: 1.037 - 1.52).
Zhu et al., (2016) RS-RAB Best prediction performance (Accuracy: 86.74%).
Bassiouni et al., (2023) Various DL approaches DL models improved accuracy by 14.46% (Softmax: 76.32%, SVM: 85.23%).
Zhu et al., (2019) RS-MultiBoosting superior performance of the RS-MultiBoosting ensemble method in forecasting
Athaudage et al., (2022) RF regression RF suitable for price prediction (MAE: 1.077, RMSE: 1.56, MAPE: 4.03%).
Liu and Huang, (2020) AdaBoosted SVM (EN-AdaPSVM) Outperforms other SVM models in noise data (Total False Rate: 11.76% (EN-AdaPSVM), 13.41% (AdaPSVM), 32.94% (CSVM)).
Zhu et al., (2017) Ensemble methods (RS-boosting and multi-boosting) RS-boosting outperforms other methods (Average Accuracy: 85.41% (RS-boosting), 84.08% (multi-boosting), 74.80% (boosting)).
Fu et al., (2022) BPNN, SVM, GA BP-GA model outperforms BP and SVM (Accuracy: BP-GA 97.19%, BP 82.33%, SVM 95.31%).
Zhang, (2022) Multimodal DL (DeepRisk) DeepRisk outperforms baseline methods.
Yang et al., (2021) Lasso-logistic model Lasso-logistic model has best prediction accuracy (Prediction Accuracy: 96.5%, Type II Error: 0.037).
Zhang et al., (2022) Fuzzy SVM (FSVM) FSVM outperform traditional SVM and DT (Higher accuracy in training and test sets).
Wang, (2021) Blockchain and fuzzy neural networks Average Accuracy: 84.11%, 84.66%(Independent Variables Set Vs).
Li, (2022) BPNN, LR BPNN and LR provide accurate predictions (Prediction Accuracy: 92.8%).
Bodendorf et al., (2023) DL model with causal inference High predictive performance and causal analysis.
Liu et al., (2022) Hybrid model (XGBoost–SMOTENC–RF) Development of a hybrid model chain, incorporating XGBoost, SMOTENC, and RF models (overall accuracy rate 91.67% and the average percentage error is 6.39%).
Rao and Li, (2022) LR, DT, integrated Logistic-RS Integrated RS improves prediction performance (Accuracy of Logistic: 0.831, Accuracy of Logistic-RS: 0.871, Accuracy of DT: 0.772, Accuracy of DT-RS: 0.795).
Handfield et al., (2020) Newsfeed analysis Predicted factory risks for a five-year planning horizon.
Pan and Miao, (2023) DL BPNN BPNN has high accuracy and low relative error (The maximum relative error calculated by the AHP is 57.41%, and the minimum is 5.88%, while the maximum relative error obtained by the BPNN is only 0.00210526%).
Wong et al., (2022) PLS-SEM, ANN AI-based models (AIRM) are significant determinants of RP and SCA.
Wang and Song, (2022) Interval Type 2 Fuzzy Neural Network (IT2FNN) IT2FNN punishes periodic deception, high accuracy (Accuracy 89%).
Nezamoddini et al., (2020) GA with NN Improved profits, lower inventory levels.
Xia et al., (2023) SVM, RF, MLP, LR RF showed best accuracy and low type I error.
Zhao and Li, (2022) SVM and BPNN Effective risk measurement, low relative error.
Li and Fu, (2022) PCA-GA-SVM PCA-GA-SVM outperforms SVM and GA-SVM, high accuracy.
Duan et al., (2021) BPNN optimized by GA Better fitting effect, higher prediction precision, and higher convergence speed.
Zhang et al., (2021) SVM optimized by firefly algorithm (FA-SVM) Improved classification efficiency and reduced error rates.
Luo et al., (2022) SVM, Particle Swarm Optimization, AdaBoost Improved classification effect, highest accuracy of 96.13%.
Dang et al., (2022) DL technology (BPNN) and blockchain Effective application of DL and blockchain.
Hosseini and Barker, (2016) Bayesian network (BN) BN allows disruption and improvement analysis.
Gao et al., (2022) Autoregressive integrated (AR), Mixture Density Networks (MDN) model, Autoregressive Integrated Moving Average (ARIMA) model, Multilayer Perceptron-Long short term memory (MLP-LSTM) model, Particle Swarm Optimization (PSO) algorithm and the improved PSO (IPSO) algorithm AR-MDN outperforms other models, IPSO has strong performance.
Wei, (2022) ML-based linear regression algorithm (ML-LRA) ML-LRA reduces supplier credit risk.
Zheng et al., (2023) Federated learning with dimensionality reduction (PCA) Collective risk prediction benefits organizations with small datasets (Maximum Difference: 0.0670 (FL vs. Local Learning), 0.0634 (FL vs. Centralized Learning).
Deiva Ganesh and Kalpana, (2022) Text-mining Use of social media data for contemporary risk understanding.
Zhang et al., (2023) Time-decayed long short-term memory (TD-LSTM), Deep Neural Network Improved predictability through TD-LSTM interpolation.
Baryannis et al., 2019a SVM and tree learning Framework offer good performance with interpretable and black-box methods (SVM Accuracy: 0.943, DT Accuracy: 0.950, Restricted DT (RDT) Accuracy: 0.916).
Han and Zhang, (2021) BPNN model BPNN effectively predicted high to extremely high risks in a food production company’s SC.
Belhadi et al., 2021b Hybrid EML approach: Gama Test (GT), Rotation Forest algorithm (RotF), Loogit Boosting algorithm (LB) Good performance of RotF-LB model (Best results for RotF-LB: Accuracy 91.72, Precision 92.96, Recall 95.53, F-measure 94.23, AUC 92.89).
Wu et al., 2022a BPNN and GA-BPNN Credit Risk Model Comparing iterations and risk prediction accuracy (GA-BPNN risk prediction accuracy above 0.92).
Lei et al., (2023) Binary image difference common area (BCAoID), SVM-SMA Highlight the effectiveness of the CGOA-SVM-SMA algorithm (Precision: 85.38%, F-score: 63%, TNR: 72%).
Yin et al., (2022) Convolutional neural network model (CNN) Optimal accuracy at 200 iterations (Comprehensive accuracy: 94.7%).
Janjua et al., (2023) Bi-LSTM CRF model Comparing Bi-LSTM CRF with Baseline CRF (the Bi-LSTM CRF model achieved an overall F1-Score of 85%, the Baseline CRF model achieved an overall F1-Score of 80%).
Yao et al., (2022) Hybrid model combining FS-MRI, SVME-AIR, and SMV AUC and KS scores for different parameters (AUC score of 0.8772 when the parameter “num_AIR” was set to 10, KS score of 0.6363 when the parameter “num_AIR” was set to 4).
Zhang et al., (2015) SVM and BPNN SVM model is superior to the BPNN model (BPNN Accuracy 66.67%, SVM model Accuracy 76.67%).
Aboutorab et al., (2022) RL-PRI approach Time efficiency of RL-PRI vs. manual process (Tarif Dispute Accuracy: 0.941463, Recession Accuracy: 0.907937, Sea Level Rise Accuracy: 0.964286, Flood Accuracy: 0.856322).
Wu, (2024) NLP techniques, word embedding models, bag-of-words approach Developed SCRisk measure, significantly correlated with operational buffers and stock return volatility.
Xu et al., (2024) Large language models (e.g., GPT-4, BERT) Improved decision support, optimized inventory management, reduced transportation costs, risk prediction.
Chen et al., (2024) GA-BPNN, LR High accuracy in predicting MSME default and credit ratings (Accuracy: 92% for default prediction, 86% for credit rating).
Zhang et al., (2025) Domain-Adaptation-based Multistage Ensemble Learning (DAMEL) DEMEL outperformed traditional models in SME credit risk evaluation, addressing data scarcity, distribution discrepancies, and class imbalance (AUC: 0.7652).
Zhou and Wang, (2025) XGBoost, RF, CNN, Least Squares SVM (LSSVM) XGBoost outperformed all models with accuracy: 91.9%, precision: 92.8%, recall: 94.7%, F1-score: 93.7%. SHAP analysis showed asset-liability ratio, quick ratio, and ESG ratings as key risk indicators.
Mojdehi et al., (2025) BallMapper + Graph Neural Network (BM-GNN) BM-GNN effectively captures financial indicators and network-based features, improving SME credit risk assessment in SC Finance.
6 Content Analysis of AI Techniques in SCRA
In recent years, the application of ML methods in SCRA has attracted substantial attention due to its potential to enhance decision-making processes and mitigate operational risks. This review paper explores various ML techniques applied in SCRA, highlighting their methodologies, key findings, and contributions.
6.1 Used ML Models and Key Findings
Table LABEL:tab:ml_methods summarizes various research methods or models used in risk prediction and analysis, along with the key findings and associated metrics from different research papers. Each row in the table represents a research paper’s first author, the research method or model used, and the specific findings and metrics achieved. These methods and findings are crucial in understanding risk prediction in various domains.
6.2 AI Models Used in SCRA
A wide range of ML and AI techniques have been applied to SC risk prediction, reflecting the increasingly data-driven nature of modern risk management. These techniques can be broadly classified into five clusters: Classical ML (CML) models, DL models, Hybrid CML+DL approaches, Ensemble methods, and Emerging Techniques such as Federated Learning, Reinforcement Learning, and Large Language Models (LLMs). Each cluster offers distinct advantages, limitations, and application contexts, with the literature providing diverse empirical evidence supporting their use. Each category is defined based on its distinct methodological characteristics, level of complexity, application scope, and suitability for specific SC risk scenarios, as shown in Figure 14.
Figure 14: Classification of AI models in SCRA based on methodology, complexity, and common application scenarios.
A detailed overview of each AI model used in SCRA, including their key characteristics, advantages, disadvantages, application cases, and suitable scenarios, is given below:
6.2.1 CML Models
CML models remain popular in SC risk prediction due to their simplicity, interpretability, and computational efficiency. Models such as Support Vector Machines (SVM), Logistic Regression (LR), Decision Trees (DT), and Bayesian Networks (BN) are particularly effective when working with structured, tabular data where relationships between variables are relatively well understood. For example, SVM and its variants have been used successfully in several studies for credit risk assessment, supplier default prediction, and operational risk evaluation (Zhu et al.,, 2016; Zhang et al.,, 2015; Zhao and Li,, 2022; Li and Fu,, 2022; Zhang et al.,, 2021; Liu and Huang,, 2020; Zhang et al.,, 2022; Fu et al.,, 2022; Liu et al.,, 2022; Xia et al.,, 2023; Lei et al.,, 2023; Gao et al.,, 2022; Yao et al.,, 2022). These models, often enhanced by feature selection or optimization algorithms, show strong classification accuracy, particularly when combined with techniques such as genetic algorithms (GA) and principal component analysis (PCA). Logistic regression remains a favored baseline due to its transparency and ease of interpretation, applied in risk assessment settings where explainability is required, such as in Li, (2022); Rao and Li, (2022); Chen et al., (2024). Decision Trees (DT), employed in studies like Baryannis et al., 2019a ; Rao and Li, (2022), provide interpretable decision rules but often suffer from overfitting unless combined into ensembles. Meanwhile, Bayesian Networks (BN) offer a probabilistic, causal perspective on risk propagation, particularly in disruption analysis settings, as demonstrated by Hosseini and Barker, (2016). Lasso regression models, which excel in feature sparsity, were applied in Yang et al., (2021) to isolate key predictors of SC risk. ML-based linear regression algorithm (ML-LRA) proposed by Wei, (2022) reduces supplier credit risk. Though highly interpretable, these models often struggle with capturing non-linear relationships and require significant domain expertise to manually engineer relevant features, limiting their scalability in high-dimensional or unstructured data environments.
Advantages
  • 1.
    Simple implementation and training.
  • 2.
    Interpretability supports regulatory compliance.
  • 3.
    Suitable for smaller datasets.
Disadvantages
  • 1.
    Limited ability to capture complex, non-linear patterns.
  • 2.
    May require extensive feature engineering.
Typical Applications
  • 1.
    Early-stage risk identification.
  • 2.
    Interpretable models for regulatory audits.
  • 3.
    Situations with limited historical data.
6.2.2 DL Models
In contrast, DL models have demonstrated superior performance in capturing complex, non-linear patterns in SC risk prediction, particularly when large, high-dimensional, or unstructured data. LSTM networks, known for their strength in sequence modeling, have been widely applied for time series forecasting in SC finance and disruption prediction (Chen et al.,, 2021; Zhang et al.,, 2023; Janjua et al.,, 2023; Gao et al.,, 2022). CNN has also been adopted, particularly when analyzing spatially structured data or feature maps, as seen in Yin et al., (2022); Zhou and Wang, (2025); Zhang et al., (2023). Multimodal DL (Zhang,, 2022; Bassiouni et al.,, 2023; Bodendorf et al.,, 2023; Pan and Miao,, 2023) models have emerged to integrate diverse data types, such as financial indicators, operational metrics, and external news sources, to enhance predictive performance, with DeepRisk showing strong results in Zhang, (2022). More recently, tensorization-based neural symbolic models (Katsaliaki et al.,, 2022) have been proposed, blending symbolic reasoning with DL for relational risk prediction. While DL models demonstrate state-of-the-art accuracy, their application in SCs is hampered by high computational costs, data-hungry training processes, and their opaque, black-box nature, which hinders regulatory acceptance in some industries.
Advantages
  • 1.
    Superior performance with complex, high-dimensional data.
  • 2.
    Effective at handling temporal and spatial dependencies.
  • 3.
    Potential to combine multimodal data.
Disadvantages
  • 1.
    High computational cost.
  • 2.
    Poor interpretability without dedicated explainability tools.
  • 3.
    Requires large labeled datasets
Typical Applications
  • 1.
    High-frequency SC data (sensor data, IoT).
  • 2.
    Multimodal data integration (text, time series, images).
  • 3.
    Long-term risk forecasting.
6.2.3 Hybrid Models (CML + DL)
Hybrid approaches combining classical ML and DL techniques aim to strike a balance between interpretability and predictive accuracy, particularly in scenarios where data characteristics vary across dimensions or stages of the risk prediction process. Several studies combined genetic algorithms (GA) with Backpropagation Neural Networks (BPNN), resulting in enhanced convergence rates and prediction precision (Wu et al., 2022a, ; Fu et al.,, 2022; Nezamoddini et al.,, 2020; Duan et al.,, 2021; Chen et al.,, 2024). In Jianying et al., (2021), PSO was also used to optimize BPNN parameters (Dang et al.,, 2022; Han and Zhang,, 2021), outperforming single models. Hybrid model chains, such as the XGBoost-SMOTENC-RF pipeline proposed in Liu et al., (2022), improve performance under class imbalance scenarios common in SC risk datasets. A particularly novel hybrid, TD-LSTM combined with deep neural networks (DNN), demonstrated strong forecasting accuracy in SC financial data in Zhang et al., (2023). By combining domain-specific feature engineering with data-driven learning, these hybrid approaches address non-linearity while preserving interpretability, though they introduce higher complexity and tuning costs.
Advantages
  • 1.
    Balance between interpretability and predictive power.
  • 2.
    Flexibility to handle different data types and distributions.
  • 3.
    Often more robust to overfitting.
Disadvantages
  • 1.
    Increased model complexity.
  • 2.
    Longer training times.
  • 3.
    Risk of suboptimal tuning in heterogeneous frameworks.
Typical Applications
  • 1.
    Medium-sized datasets with mixed data types.
  • 2.
    Cases requiring both prediction accuracy and interpretability.
  • 3.
    Data with substantial noise and missing values.
6.2.4 Ensemble Models (Bagging, Boosting, Stacking)
Ensemble methods, particularly RF (Athaudage et al.,, 2022; Wang et al.,, 2022, 2021; Xia et al.,, 2023; Liu et al.,, 2022; Zhu et al.,, 2019), boosting algorithms (XGBoost, AdaBoost) (Liu et al.,, 2022; Zhou and Wang,, 2025; Zhang et al.,, 2025), and multi-boosting approaches consistently outperform individual SC risk prediction models by aggregating predictions from multiple weak learners. Studies such as Zhu et al., (2019, 2017); Liu and Huang, (2020); Belhadi et al., 2021b demonstrated that multi-boosting ensembles effectively handle data noise and feature redundancy, improving accuracy in risk classification tasks. Cost-sensitive Random Forest (CSL-RF) models further enhance robustness under unbalanced risk categories, as shown in Wang et al., (2022). Novel approaches like the Domain-Adaptation-based Multistage Ensemble Learning (DAMEL) framework, proposed in Zhang et al., (2025), directly address data scarcity, distribution shifts, and class imbalance, particularly relevant for SME credit risk prediction. While ensembles excel in predictive accuracy, their major trade-off is reduced interpretability, especially as ensemble size increases.
Advantages
  • 1.
    Enhanced predictive performance.
  • 2.
    More robust to overfitting.
  • 3.
    Can combine strengths of multiple algorithms.
Disadvantages
  • 1.
    Computationally expensive.
  • 2.
    Poor interpretability.
  • 3.
    Difficult to tune optimally.
Typical Applications
  • 1.
    Large, noisy datasets.
  • 2.
    Situations where prediction accuracy is paramount.
  • 3.
    Multi-source data environments.
6.2.5 Emerging and Hybrid Techniques
Several cutting-edge techniques have emerged, demonstrating strong potential for real-time risk monitoring and collaborative prediction across SC partners. Federated learning (FL) approaches, such as those in Zheng et al., (2023), enable decentralized risk modeling across firms without exposing sensitive data, a crucial advantage in privacy-sensitive SC networks. Reinforcement Learning (RL), employed in Aboutorab et al., (2022), optimizes adaptive intervention strategies, balancing risk mitigation costs and performance outcomes. Graph Neural Networks (GNN), such as the BM-GNN framework introduced in Mojdehi et al., (2025), excel at capturing relational dependencies between suppliers, enriching risk propagation modeling. Lastly, natural language processing (NLP) and LLMs, including GPT-4 and BERT, have been applied for text-based risk identification, extracting risk signals from news, social media, and operational reports (Wu,, 2024; Xu et al.,, 2024; Deiva Ganesh and Kalpana,, 2022).
Advantages
  • 1.
    Enables collaborative modeling across multiple organizations.
  • 2.
    Preserves data privacy.
  • 3.
    Hybrid symbolic-ML interpretability.
Disadvantages
  • 1.
    Requires careful coordination across participants.
  • 2.
    Vulnerable to communication bottlenecks.
  • 3.
    Model updates may suffer from drift across sites.
Typical Applications
  • 1.
    Collaborative supplier monitoring.
  • 2.
    Risk prediction across supply networks.
  • 3.
    Situations where data privacy regulations restrict data sharing.
7 Practical and Managerial Implications
The findings of this article will play a significant role in managerial strategies designed to enhance SCRM. First, it stresses that ML and AI are important tools that will give rise to a new phase of SCRA. Managers are encouraged to try advanced ML models, such as RF, XGBoost, and blends. They stimulate considerable risk assessment precision. These models are likely to enhance the accuracy of risk evaluation enormously and, thus, the effectiveness of their risk management strategies.
In addition, the study highlights the importance of flexibility and adaptability in an era after COVID-19, suggesting that companies need new SC strategies (Sheng et al.,, 2021). Managers must now place contingency plans at the forefront, focusing on those not readily broken by future disruptions. It is only by doing routine work thoroughly that one can concentrate on something special without fear or disturbance from the outside world, using people who have already developed a variety of fresh paths, so avoiding crisis after crisis with all that effort having been wasted.
The findings also highlight the importance of ensemble methods like RF and XGBoost in risk assessment (Zhou and Wang,, 2025). It is possible that such models could improve the accuracy of risk prediction and may be of interest to managers striving to strengthen their SCRM systems. The importance of time series prediction models, such as ARIMA and LSTM, cannot be overstated. Companies are advised to create tools predicting potential risks based on historical performance to approach the task of SC disruption mitigation with much-needed foresight.
ML models must produce precise and human-readable outputs as they assist decision-makers in understanding the process of risk assessment and even its findings. Furthermore, the study indicates the significance of real-time data mining and using non-traditional sources like social media, e.g., Twitter, as critical components of full-scale risk identification (Deiva Ganesh and Kalpana,, 2022). Managers should be careful and up-to-date in understanding real-time information.
Blockchain, combined with ML models, represents a powerful tool for enhancing transparency, traceability, and risk monitoring in the SC. Blockchain ensures secure and immutable data sharing, while ML models enable predictive risk assessment and anomaly detection, making risk management more proactive and efficient (Dang et al.,, 2022). Organizations exploring blockchain solutions should assess how integrating ML-driven risk forecasting can enhance SC resilience and optimize risk mitigation strategies. SMEs require principal strategies in risk evaluation regarding credit risks since they have distinct characteristics and risks in terms of operations. Conversely, implementing strategies based on SMEs’ needs and constraints is required. Moreover, federated learning should be implemented when data privacy issues arise. This method addresses the risk but protects confidential information, permitting partners to conduct joint risk forecasting. Compliance with security and data privacy must be performed almost immediately.
Several hybrid risk assessment methodologies that combine various ML models and techniques bring the necessary comprehensive view of SCRA (Mojdehi et al.,, 2025). These approaches enable managers to explore advanced risk management strategies, enhancing their ability to effectively identify, assess, and mitigate risks. Additionally, comparing benchmarks of different ML models is critical to understanding more information about the best ones for particular risk scenarios. The overall message of the study is outlining the need for continuous performance evaluations and improvements. Risk assessment strategies should be able to evolve alongside other circumstances. Other assumptions refer to integrating ML-based risk assessment into already existing SC workflows and decision-making patterns. It is necessary to determine how such technologies can be easily implemented. Employee training in data analysis, ML, and AI is critical. Managers should invest in skill development to empower their SC professionals to harness these technologies effectively. Collaboration and knowledge sharing within organizations and across industries are encouraged to collectively build robust risk management strategies. Such partnerships have the potential to enhance risk resilience significantly.
Lastly, understanding and adhering to data protection and privacy regulations is critical when deploying AI and ML for risk assessment. Managers should ensure that their risk management practices align with the legal requirements to avoid compliance issues.
8 Challenges and Limitations
Integrating AI in SCRA brings forth several challenges and limitations that researchers and practitioners must address. These challenges encompass the following:
  1. 1.
    Data Availability and Quality: AI models in SCRA rely heavily on data. However, ensuring the availability and quality of data remains a persistent challenge. SC data often comes from various sources, which can be fragmented, incomplete, or outdated (Schöpper and Kersten,, 2021). Researchers must explore ways to consolidate, cleanse, and augment data to enhance the effectiveness of AI models.
  2. 2.
    Interpretability and Explainability: Many AI techniques, such as DL and neural networks, are often considered “black-box” models. This lack of interpretability and explainability is a significant limitation (Singh et al.,, 2024). In SCRA, stakeholders require insights into why a model makes a particular prediction or decision. Developing AI models that are not only accurate but also interpretable is crucial for their adoption in real-world SCM.
  3. 3.
    Dynamic and Evolving Risk Factors: The SC environment is dynamic and continuously evolving (Nitsche et al.,, 2021). New risk factors emerge, and existing ones change over time. AI models may struggle to adapt to these dynamic conditions. Researchers need to explore techniques for real-time risk assessment and adaptive models that can respond to changing risk profiles.
  4. 4.
    Integration with Existing Systems: Integrating AI solutions into existing SCM systems can be challenging due to various drivers (technological, organizational, environmental, and human) and barriers (technical, organizational, economic, and human) that influence adoption (Shahzadi et al.,, 2024). Organizations often rely on legacy systems, and ensuring compatibility and seamless integration can pose hurdles. Research in this area should focus on creating modular AI solutions that can be integrated into diverse systems with minimal disruption.
9 Future Research Directions
In light of the challenges and limitations, future research directions in AI-based SCRA are multi-faceted and have the potential to advance the state-of-the-art. The development of AI models that deliver accurate predictions and offer comprehensive transparency and interoperability should be emphasized. Techniques like Local Interpretable Model-Agnostic Explanations (LIME) and SHapley Additive exPlanations (SHAP) serve as pivotal approaches to enhance the interpretability of AI models (Olan et al.,, 2024). By employing these methods, the goal is to demystify the decision-making process of AI models, enabling stakeholders to comprehend the rationale behind specific predictions or decisions. This transparency enhances trust and facilitates more effective integration of AI-driven insights into SC risk management strategies.
Combining multiple AI techniques, such as ML and expert systems, can enhance the robustness of SCRA. Future research should explore hybrid AI models that leverage the strengths of different approaches to improve prediction accuracy and reliability (Mojdehi et al.,, 2025). Addressing data quality challenges involves multifaceted strategies. Utilizing larger datasets for analysis, as emphasized by (Wang et al.,, 2021), enhances the robustness of risk assessment, enabling a more comprehensive evaluation of risk factors and improving predictive accuracy. Additionally, initiatives should be aimed at democratizing access to standardized SC data repositories. Creating and maintaining such repositories enables stakeholders across the SC spectrum to access reliable and consistent data. Increased accessibility fosters better-informed decision-making processes and promotes using standardized data sets for improved AI model training and validation. Investigating AI models capable of real-time risk assessment will be pivotal. These models should adapt to emerging risks and provide timely alerts or recommendations for risk mitigation.
Research on creating integration frameworks for AI-based SCRA is essential (Shahzadi et al.,, 2024). These frameworks should facilitate easy integration into existing SC management systems, ensuring a smooth transition to AI-driven risk assessment. Expanding the application of AI in risk assessment to various domains within the SC, including logistics, procurement, and production, offers opportunities for comprehensive risk management. Investigate AI models capable of adapting to pandemic-induced disruptions within SCs. Focus on understanding the unique dynamics and challenges posed by COVID-19, aiming to develop AI-driven strategies that enable resilience and agility in such crisis scenarios. As AI plays an increasingly significant role in decision-making, researchers should also address ethical and legal considerations. This includes issues related to data privacy, bias, and fairness in AI models. By addressing these challenges and pursuing these research directions, AI-based SCRA can make significant strides toward more effective and reliable risk management in an ever-evolving global SC landscape.
10 Conclusions
This extensive review delved into the SCRA domain and examined the complex challenges of integrating AI, especially ML. Our thorough analysis encompassed a substantial pool of 1,903 SCRA papers, ultimately narrowing our focus to a refined subset of 54 papers. This endeavor has yielded significant insights and key findings, making substantial contributions to the SCRA literature.
Our primary contributions can be summarized by addressing the four key research questions:
  • RQ-1:
    Our review thoroughly analyzes the existing state of research concerning the application of AI and ML techniques in SCRM. We have meticulously assessed a substantial body of literature to extract meaningful insights, specifically from 54 selected articles.
  • RQ-2:
    Through our analysis, we have identified the AI and ML techniques most commonly utilized in the domain of SCRA. This clarifies the methodologies employed and underscores the trends shaping the field.
  • RQ-3:
    Our review highlights the key findings and trends prevalent in the literature, shedding light on the progress, developments, and major areas of focus within AI and ML in SCRM.
  • RQ-4:
    We have identified critical research gaps and proposed future directions to guide further explorations in this field, ensuring the continued evolution of SCRA methods.
Despite these contributions, our research has some limitations. Our data primarily originated from Google Scholar and WoS rather than Scopus, potentially introducing biases. Our focus on papers published between 2015 and 2025 presents a limitation, as it may omit earlier relevant research that could provide valuable historical context. Additionally, our restriction to English-language papers may have omitted valuable non-English research.
Looking forward, several promising avenues for future research emerge in the domain of SCRA using AI. These directions encompass operationalizing interconnectedness, transformability, and sharing within SCRA frameworks, investigating evolving ICT roles in prediction and response, revisiting the cost implications of resilience enhancement, and exploring emerging technologies like blockchain. These endeavors are expected to enhance risk assessment effectiveness in the post-COVID era. Reviewers should adopt a comprehensive approach to gain deeper insights into this field by expanding their search to diverse databases, including non-English and grey literature sources, using snowballing techniques, collaborating with experts from related disciplines, leveraging ML tools for data analysis, and staying up-to-date with the latest research. Combining various search methods and expert opinions, such a multidisciplinary strategy is essential for uncovering valuable insights and emerging trends in this dynamic and critical field.
Our SLR, driven by these four primary research questions, has contributed valuable insights into the AI/ML application in SCRA, identified common techniques, outlined key findings and trends, and proposed essential research directions. This work serves as a guide for both researchers and practitioners, facilitating advancements in SCRM.